OpenStudy (anonymous):

The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression 80 *2^t/28 grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression represent?

OpenStudy (anonymous):

\[80\times 2^{\frac{0}{28}}=80\times 2^0=80\times 1=80\] for the first one others are similar

OpenStudy (anonymous):

Thank you (: